# Nonlocal Models in Biology and Life Sciences: Sources, Developments, and Applications

### v1

@@ Abstract
Mathematical modelling is one of the fundamental techniques for understanding biophysical mechanisms in developmental biology. It helps researchers to analyze complex physiological processes and connects like a bridge between theoretical and experimental observations. Various groups of mathematical models have been studied to analyze these processes, and the nonlocal models are one of them. Nonlocality is important in realistic mathematical models of physical and biological systems at small-length scales. It characterizes the properties of two individuals located in different locations. This review illustrates different nonlocal mathematical models applied to biology and life sciences. The major focus has been given to sources, developments, and applications of such models. Among other things, a systematic discussion has been provided for the conditions of pattern formations in biological systems of population dynamics. Special attention has also been given to nonlocal interactions on networks, network coupling and integration, including models for brain dynamics that provide us with an important tool to better understand neurodegenerative diseases. In addition, we have discussed nonlocal modelling approaches for cancer stem cells and tumor cells that are widely applied in the cell migration processes, growth, and avascular tumors in any organ. Furthermore, the discussed nonlocal continuum models can go sufficiently smaller scales applied to nanotechnology to build biosensors to sense biomaterial and its concentration. Piezoelectric and other smart materials are among them, and these devices are becoming increasingly important in the digital and physical world that is intrinsically interconnected with biological systems. Additionally, we have reviewed a nonlocal theory of peridynamics, which deals with continuous and discrete media and applies to model the relationship between fracture and healing in cortical bone, tissue growth and shrinkage, and other areas increasingly important in biomedical and bioengineering applications. Finally, we provided a comprehensive summary of emerging trends and highlighted future directions in this rapidly expanding field.
@@
Nonlocal models; nonlocal interactions in time and space; cell biology, genomics, and populations dynamics; network coupling and integration; epidemiology and immunology; pattern formations; nonequilibrium phenomena and processes; health sciences and innovative technologies; smart and intelligent systems; active matter and AI.
@@ 1 Introduction
Nonlocal interactions are an intrinsic feature of many natural processes and phenomena studied in life sciences. Mathematical models assist such studies with a rigorous foundation and, together with computational models and techniques, provide additional vital insights, complementing experimental investigations. In 1923, Dr. W. M. Feldman introduced the term “biomathematics” in the book title, a highly relevant discipline nowadays feldman1923 . He was a physician and was more interested in the numerical data of some of his patients’ more common dynamics. In 1938, a theoretical physicist Dr. N. Rashevsky, published the first book on mathematical biology and mathematical biophysics entitled “Mathematical Biophysics: Physico-Mathematical Foundations of Biology” rashevsky1938 . Due to many technological advances since then, biomathematics has evolved from Feldman’s simple “mathematical principles for biology students” to one of the most promising tools for biology and medicine. Mathematical models help break down their dynamic processes into individual elements to study a natural biological process. Moreover, these models can bridge the gap between proposed molecular interactions inside and between cells and their tissue-level effects. As they developed, the hybrid field of “biostatistics” emerged and helped to analyze different scientific issues mathematically, e.g., biodiversity, agriculture, medicine, etc. Alan Turing also contributed to biomathematics by discovering the condition for particular shapes of each organism, and he is known as “the founder of contemporary Mathematical Biology”. In addition, his works included three main ingredients: modelling, differential equations, and numerical solutions amturing .
@@
Mathematical models deal with one or more equations (discrete-time, continuous- and fractional-time ordinary and partial differential equations) depending on the component(s) of the system jdmurray ; edelstein2005 ; tomlin2007 . In addition, one or more parameters may be involved in the model to capture its different characteristics vandermeer2010 ; mittelbach2019 . Discrete-time models governed by difference equations are more appropriate for organisms with non-overlapping generations, e.g., an insect population with one generation per year azuaje2011 ; azizi2021 . On the other hand, ordinary and/or partial differential equations are used for continuous-time models and are appropriate for organisms with overlapping generations cappuccino1995 . In addition, fractional-time models are mainly used to incorporate the memory effect of the organisms in time. In this article, we cover both continuous- and fractional-time models.
@@
The model described by an ordinary differential equation (ODE) queries the state of an entity at a given time; hence, only the time derivatives are present in such systems. For instance, the simplest SIR model in epidemiology helps to predict different things, such as how a disease spreads, the total number infected, and the duration of an epidemic ross1916 ; ross1917 ; kermack1927 . These models also deal with cell signalling pathways, population growth, enzymatic inhibitor reactions, and ecological models. In studying the temporal dynamics of a system, the major concern is shifting one equilibrium to another equilibrium or non-equilibrium (e.g., limit cycle) by a small perturbation around it jdmurray ; jones2009 ; perko2013 . The temporal dynamics of a model are described by its reaction terms, which mainly include birth, death, and interactions between individuals. Sometimes, a small perturbation may completely change the dynamics of a system, known as butterfly effects in chaos theory lorenz1963 . Along with the ODE models, fractional-order differential equation (FODE) models also play a crucial role in biological systems. The fractional differentials are interpreted in terms of Riemann-Liouville miller1993 , Caputo miller1993 , Caputo-Fabrizio caputo2015 , and Riesz cheng2019 fractional derivatives. FODE provides a powerful tool for describing the memory of different substances and the nature of the inheritance, and memory and genetic properties are the two key factors in biological modelling, e.g., membranes of cells have fractional order electrical conductance cole1933 , cell rheological behaviour djordjevic2003 . FODE is also studied in the epidemic model of SIR type to understand the memory effect on disease propagation saeedian2017 .
@@
The use of spatio-temporal models has increased in the scientific literature, and it has additional benefits over temporal studies auchincloss2012 ; yu2018 ; byun2021 . It can capture localized clusters or heterogeneity of the space-time interactions between the components of a system. The word heterogeneity describes the difference or diversity in a system, which means “lots of different things”. Such differences can occur in space and time, and in ecology, this term is mainly used to describe the land cover or habitat fryxell2014 . On the other hand, a homogeneous system is a single-phase system of continuous variables, and each of these variables is distributed uniformly throughout the domain of definition. This homogeneity or heterogeneity depends on the scale considered salevin . For example, one heterogeneous habitat can be seen fifty meters from its top, but it can look like a homogeneous habitat from five hundred meters or above. Looking at the blood with a bare eye seems homogeneous, but it has different components like red blood cells, plasma, and platelets under the microscope. In reality, heterogeneity comes when every member has different values for the characteristics of our interests. Well-known reaction-diffusion (RD) equations can capture this type of heterogeneous phenomenon. These equations are conventionally used in chemical interactions to describe the concentrations of the chemicals at different times in a system but are later applied to different fields page2003pattern ; colizza2007reaction ; song2019spatial ; wuyts2019tropical ; kholodenko2003 . The Gierer-Meinhardt model and its relevant modifications capture different biological processes such as animal skin patterns, nerve cells and their activities, organ formation, etc. meinhardt1982 ; meinhardt2009 . In chemistry, the well-known CIMA reaction can generate different types of spatial patterns, explained by RD modelling castets1990 ; de1991 ; lengyel1991 . These spatio-temporal models also describe health outcomes, such as contracting a disease, in different locations and spatial points in time diggle2019 .
@@
Starting with Fisher’s (1937) seminal work on gene spread, researchers have developed various ways of modelling spatial interactions in biology. Nowadays, nonlocality is emphasised in mathematical biology. This manuscript focuses on different nonlocal models based on individual interactions, which may help biologists capture essential features of spatial phenomena. The concept of nonlocality exists everywhere, and it is essential in realistic mathematical models of many aspects of the physical world silling2019 . In an ecosystem, organisms compete for resources to survive, grow, and reproduce. Animals compete for air, food, shelter, water, and space. Plants compete for needed resources, including air, water, sunlight, and space. Therefore, in ecology, there is no real justification for assuming the interactions as local since (i) the population always competes for shared resources, (ii) the individuals in a population communicate either visually or through chemical means furter1989 . In addition, the species sense their surroundings for making different decisions, e.g., the predator finds the prey or the prey avoids the predator eftimie2007 ; levine2000 . High population density at a spatial location impacts the same spatial point and neighbouring locations, so considering the impact of long-range effects is important. Many particles sense signals over extended regions, e.g., filopodia in cells detect signals at multiple cell diameters away chen2020 ; kornberg2014 . Along with these factors, dispersal distances may also be nonlocal, e.g., seeds can be transported significantly from the source. The spatial convolution is more realistic than the usual diffusion for many ecological populations sherratt2016 . Different methods have been used for kernel estimation, including dispersive individuals nathan2012 .
@@
The usual RD model in interacting species can capture only the spatial interactions of neighbourhood locations, generally called short-range interactions. On the other hand, the nonlocal term incorporates the behaviour of a distribution function at several points, and a system of integro-differential equations describes it. Generally, these nonlocal models deal with long-range interactions in the interacting populations. Along with these nonlocal RD models, nonlocal advection-RD models have also gained more attention in recent years. These models have the capacity to include cell-cell adhesion interactions for tissue dynamics chen2020 . Nonlocal mathematical models are also used to investigate the progress of an epidemic, including age-structured modelling kendall1965 ; kang2021 . Furthermore, a system of integro-differential equations is used to describe the general nonlocal continuum theory, which helps to make bridge-base sensors and is applied to biological sciences eringen2003 ; lavrik2004 ; ekinci2005 ; lu2006 . In addition, the spatial dispersal mechanism of cells or organisms is one of the central topics in biology and ecology. Random dispersal has been used to describe the movement of organisms between adjacent spatial locations. However, for some organisms, the movements can occur between non-adjacent spatial locations, and the nonlocal dispersal captures such behaviours hutson2003 .
@@
Multiscale modelling approaches of materials can capture one or more physical phenomena at different lengths and/or time scales nieminen2002 ; shaat2020 . Materials generally exhibit and secrete different phenomena at different lengths and time scales shaat2020 . Due to the lack of measures and length scales in classical mechanics, the material’s microstructural features (e.g., micro-slipping and twinning), geometries, and/or topologies (e.g., band-gap) are unpredictable. Hierarchical and hybrid are the two types of multiscale modelling approaches studied in the literature. In the case of the hierarchical multiscale model, the material is broken down into multiple scales, and models at the different scales are run independently. In the end, coupling parameters are introduced to couple those scales. On the other hand, all different scales work together as a single package for hybrid multiscale models. The material consists of interconnected multiple subdomains over short and long ranges, and the nonlocal continuum mechanics model can capture their mechanics.
@@
Nonlocal models are capable of capturing some behaviour of biological materials such as cortical bone fracture deng2009 , the rupture in lipid membranes taylor2016 , tumor shrinkage lejeune2021 , etc. Peridynamics is one of the main theoretical and computational frameworks that unify the mechanics of discrete and continuous media lejeune2021 . Since the first papers were published, peridynamics methodology has been used to model discontinuous and long-range forces using pair-wise interactions between particles silling2000 . This methodology is nonlocal due to considering the long-range interaction between particles. As it deals with continuous and discrete media, it is a compelling model of biological tissue on the cell population scale. In addition, piezoelectric and other smart materials are the most commonly used medical materials. For instance, ultrasound scanners use a piezoelectric transducer; they use the converse piezoelectric effect to send out sound vibrations reflected by muscles, organs, etc. manjon2018 .
@@
Mathematical and computational models play a pivotal role in unravelling the intricate and dynamic processes within the realm of life sciences. These models provide a systematic framework to analyze, simulate, and predict biological phenomena, aiding researchers and practitioners in understanding the underlying principles governing life at various levels of complexity. Life sciences encompass a vast array of disciplines, from molecular biology and genetics to ecology and neuroscience. Understanding the behaviour and interactions of biological systems requires the integration of data and theories into mathematical and computational frameworks. These models enable the abstraction of biological processes into mathematical equations, algorithms, or simulations, allowing for a quantitative and often more precise representation of biological phenomena.
@@ 2.1 Homogeneous and heterogeneous systems
Homogeneous and heterogeneous systems are fundamental concepts in mathematical biology that play crucial roles in understanding various biological phenomena. Homogeneous systems are the baseline for understanding the fundamental dynamics of biological systems before incorporating more complex factors. For instance, patterns are observed in nature and have visible regularities. These patterns recur in different contexts and can sometimes be modelled mathematically. According to Turing amturing , “when two chemical species with different diffusion rates react with each other, the spatially homogeneous state may become unstable, thereby leading to a nontrivial spatial structure”. This looks counter-intuitive as it is expected to be a uniform distribution of the chemical substances through the diffusion term. In general, the diffusion term in RD equations contributes to transferring the concentrations throughout the spatial domain, and the linear or nonlinear reaction terms come from the mass action laws jdmurray ; volpert2009 . Without loss of generality, we can consider a general form of RD equations with $m$-substances $\mathbf{u}(\mathbf{x},t)=(u_{1}(\mathbf{x},t),\ldots,u_{m}(\mathbf{x},t))^{T}$ at time $t>0$, and at $\mathbf{x}$ in a spatial domain $\Omega\subseteq\mathbb{R}^{n}$ as jdmurray :
@@
Depending on the parameters and interactions between organisms, the RD equation (1) exhibits different types of dynamic non-homogeneous solutions, such as travelling waves, wave-trains, spatio-temporal chaos, etc. In addition, reaction-diffusion systems can produce non-homogeneous stationary patterns such as spots, stripes, and target patterns jdmurray . In 1937, Fisher and Kolmogorov, Petrovskii, and Piskunov introduced the scalar RD equation kolmogorov1937 :
@@
and studied the existence of the travelling wave solutions along with the propagation of the wave speed. The equation (2) is known as the Fisher-KPP equation. Researchers have been studying and applying this Fisher-KPP model in various fields, e.g., physics, chemistry, biology, and medicine. A particular nonlinear form $g(u)=u(1-u)$ of the RD equation (2) has been taken a lot of attention in capturing different phenomena through mathematical modelling, e.g., spatial spreading of cell populations simpson2013 , wound-healing cell migration sherratt1990 ; maini2004 , protein misfolding fornari2019 ; schafer2021 , biological invasion petrovskii2005 ; skellam1951 ; levin2003 , etc.
As discussed earlier, the RD equations (1) are capable of producing different types of stationary and non-stationary non-homogeneous patterns. Several mechanisms are available for getting non-stationary patterns jdmurray . Turing suggested a mechanism to obtain some of the non-homogeneous stationary patterns mathematically, called the Turing pattern amturing . One of the main conditions for existing Turing patterns requires at least two interacting populations for the RD equations. Therefore, the usual Turing mechanism can not be applied to the single-species local RD model (2). However, it can be applied to a single species RD model with nonlocal interactions. We first describe a nonlocal version of the Fisher-KPP RD model for a single species population (2), then move to the multi-species populations and other nonlocal interactions.

### v2

@@ Abstract
Mathematical modelling is one of the fundamental techniques for understanding biophysical mechanisms in developmental biology. It helps researchers to analyze complex physiological processes and connects like a bridge between theoretical and experimental observations. Various groups of mathematical models have been studied to analyze these processes, and the nonlocal models are one of them. Nonlocality is important in realistic mathematical models of physical and biological systems when local models fail to capture the essential dynamics and interactions that occur over a range of distances (e.g., cell-cell, cell–tissue adhesions, neural networks, the spread of diseases, intra-specific competition, nanobeams, etc.). This review illustrates different nonlocal mathematical models applied to biology and life sciences. The major focus has been given to sources, developments, and applications of such models. Among other things, a systematic discussion has been provided for the conditions of pattern formations in biological systems of population dynamics. Special attention has also been given to nonlocal interactions on networks, network coupling and integration, including brain dynamics models that provide an important tool to understand neurodegenerative diseases better. In addition, we have discussed nonlocal modelling approaches for cancer stem cells and tumor cells that are widely applied in the cell migration processes, growth, and avascular tumors in any organ. Furthermore, the discussed nonlocal continuum models can go sufficiently smaller scales, including nanotechnology, where classical local models often fail to capture the complexities of nanoscale interactions, applied to build biosensors to sense biomaterial and its concentration. Piezoelectric and other smart materials are among them, and these devices are becoming increasingly important in the digital and physical world that is intrinsically interconnected with biological systems. Additionally, we have reviewed a nonlocal theory of peridynamics, which deals with continuous and discrete media and applies to model the relationship between fracture and healing in cortical bone, tissue growth and shrinkage, and other areas increasingly important in biomedical and bioengineering applications. Finally, we provided a comprehensive summary of emerging trends and highlighted future directions in this rapidly expanding field.
@@
Nonlocal models; nonlocal interactions in time and space; nonlocal processes and phenomena; cell biology, genomics, and populations dynamics; network coupling and integration; epidemiology and immunology; pattern formations; nonequilibrium phenomena and processes; health sciences and innovative technologies; smart and intelligent systems; active matter and AI
@@ 1 Introduction
Nonlocal interactions are among the fundamental aspects of many natural processes and phenomena studied in the life sciences, e.g., cell-cell communication, brain function, disease transmission, ecological population dynamics, etc. Mathematical models assist such studies with a rigorous foundation and, together with computational models and techniques, provide additional vital insights, complementing experimental investigations. In 1923, Dr. W. M. Feldman introduced the term “biomathematics” in the book title, a highly relevant discipline nowadays [1]. He was a physician and was more interested in the numerical data of some of his patients’ more common dynamics. In 1938, a theoretical physicist Dr. N. Rashevsky, published the first book on mathematical biology and mathematical biophysics entitled “Mathematical Biophysics: Physico-Mathematical Foundations of Biology” [2]. Due to many technological advances since then, biomathematics has evolved from Feldman’s simple “mathematical principles for biology students” to one of the most promising tools for biology and medicine. Mathematical models help break down their dynamic processes into individual elements to study a natural biological process. Moreover, these models can bridge the gap between proposed molecular interactions inside and between cells and their tissue-level effects. As they developed, the hybrid field of “biostatistics” emerged and helped to analyze different scientific issues mathematically, e.g., biodiversity, agriculture, medicine, etc. Alan Turing also contributed to biomathematics by discovering the condition for particular shapes of each organism, and he is known as “the founder of contemporary Mathematical Biology”. In addition, his works included three main ingredients: modelling, differential equations, and numerical solutions [3].
@@
Mathematical models are used to understand, predict, and analyze complex biological systems and phenomena by translating biological processes into mathematical terms. They deal with one or more equations (continuous- and fractional-time ordinary and partial differential equations) depending on the component(s) of the system [4, 5, 6, 7]. In addition, one or more parameters may be involved in the model to capture its different characteristics [8, 9]. Discrete-time models governed by difference equations are more appropriate for organisms with non-overlapping generations, e.g., an insect population with one generation per year [10, 11]. On the other hand, ordinary and partial differential equations are used for continuous-time models and are appropriate for organisms with overlapping generations [12]. In addition, fractional-time models are mainly used to incorporate the memory effect of the organisms in time. In this article, we cover both continuous- and fractional-time models.
@@
The model described by an ordinary differential equation (ODE) queries the state of an entity at a given time; hence, only the time derivatives are present in such systems. For instance, the simplest SIR model in epidemiology helps to predict different things, such as how a disease spreads, the total number infected, and the duration of an epidemic [13, 14, 15]. These models also deal with cell signalling pathways, population growth, enzymatic inhibitor reactions, and ecological models. In studying the temporal dynamics of a system, the major concern is shifting one equilibrium to another equilibrium or non-equilibrium (e.g., limit cycle) by a small perturbation around it [4, 16, 17]. The temporal dynamics of a model are described by its reaction terms, which mainly include birth, death, and interactions between individuals. Sometimes, a small perturbation may completely change the dynamics of a system, known as butterfly effects in chaos theory [18]. Along with the ODE models, fractional-order differential equation (FODE) models also play a crucial role in biological systems. The fractional differentials are interpreted in terms of Riemann-Liouville [19], Caputo [19], Caputo-Fabrizio [20], and Riesz [21] fractional derivatives. FODE provides a powerful tool for describing the memory of different substances and the nature of the inheritance, and memory and genetic properties are the two key factors in biological modelling, e.g., membranes of cells have fractional order electrical conductance [22], cell rheological behaviour [23]. FODE is also studied in the epidemic model of SIR type to understand the memory effect on disease propagation [24].
@@
The use of spatio-temporal models has increased in the scientific literature, and it has additional benefits over temporal studies [25, 26, 27]. It can capture localized clusters or heterogeneity of the space-time interactions between the components of a system. The word heterogeneity describes the difference or diversity in a system, which means “lots of different things”. Such differences can occur in space and time, and in ecology, this term is mainly used to describe the land cover or habitat [28]. On the other hand, a homogeneous system is a single-phase system of continuous variables, and each of these variables is distributed uniformly throughout the domain of definition. This homogeneity or heterogeneity depends on the scale considered [29]. For example, one heterogeneous habitat can be seen fifty meters from its top, but it can look like a homogeneous habitat from five hundred meters or above. Looking at the blood with a bare eye seems homogeneous, but it has different components like red blood cells, plasma, and platelets under the microscope. In reality, heterogeneity comes when every member has different values for the characteristics of our interests. Well-known reaction-diffusion (RD) equations can capture this type of heterogeneous phenomenon. These equations are conventionally used in chemical interactions to describe the concentrations of the chemicals at different times in a system but are later applied to different fields [30, 31, 32, 33, 34]. The Gierer-Meinhardt model and its relevant modifications capture different biological processes such as animal skin patterns, nerve cells and their activities, organ formation, etc. [35, 36]. In chemistry, the well-known CIMA reaction can generate different types of spatial patterns, explained by RD modelling [37, 38, 39]. These spatio-temporal models also describe health outcomes, such as contracting a disease, in different locations and spatial points in time [40].
@@
Starting with Fisher’s (1937) seminal work on gene spread, researchers have developed various ways of modelling spatial interactions in biology. This manuscript focuses on different nonlocal models based on individual interactions, which can help biologists capture essential features of spatial phenomena. The concept of nonlocality exists everywhere, and it is essential in realistic mathematical models of many aspects of the physical world [41, 42]. These models are useful for capturing spatially distributed phenomena in biological systems, such as movement, diffusion, or interaction among organisms, cells, or chemicals not confined to adjacent interactions about their positions. Several key factors are involved when analyzing nonlocal models in biology. For instance, the strength of interactions between two points is determined by how interactions occur over a distance as a function of the distance between them. Many biological systems have boundaries that affect behaviour at a distance, such as skin or organ membranes, influencing diffusion and interaction [43]. In epidemiology, nonlocal models are employed to simulate the transmission of diseases over extensive geographical areas, where persons from far locations may come into contact, hence spreading illnesses through travel or migration [44]. Neuronal nonlocal models in neuroscience represent the interactions between physically distant neurons yet functionally linked by long-range synaptic connections [45, 46]. In an ecosystem, organisms compete for resources to survive, grow, and reproduce. Animals compete for food, shelter, water, and space, while plants compete for resources, including air, water, sunlight, and space. Therefore, in ecology, there is no real justification for assuming the interactions as local since (i) the population always competes for shared resources, (ii) the individuals in a population communicate either visually or through chemical means [47]. In addition, the species sense their surroundings for making different decisions, e.g., the predator finds the prey or the prey avoids the predator [48, 49, 50, 51]. High population density at a spatial location impacts the same spatial point and neighbouring locations, so considering the impact of long-range effects is important. Many particles sense signals over extended regions, e.g., filopodia in cells detect signals at multiple cell diameters away [52, 53]. Along with these factors, dispersal distances may also be nonlocal, e.g., seeds can be transported significantly from the source. The spatial convolution is more realistic than the usual diffusion for many ecological populations [54]. Different methods have been used for kernel estimation, including dispersive individuals [55].
@@
The usual RD model in interacting species can capture only the spatial interactions of neighbourhood locations, generally called short-range interactions. On the other hand, the nonlocal term incorporates the behaviour of a distribution function at several points, and a system of integro-differential equations describes it. Generally, these nonlocal models deal with long-range interactions in the interacting populations. Along with these nonlocal RD models, nonlocal advection-RD models have also gained more attention in recent years. These models can include cell-cell adhesion interactions for tissue dynamics [52, 56]. Nonlocal mathematical models are also used to investigate the progress of an epidemic, including age-structured models [57, 58]. Furthermore, a system of integro-differential equations is used to describe the general nonlocal continuum theory, which helps to make bridge-base sensors and is applied to biological sciences [59, 60, 61, 62]. In addition, the spatial dispersal mechanism of cells or organisms is one of the central topics in biology and ecology. Random dispersal has been used to describe the movement of organisms between adjacent spatial locations. However, for some organisms, the movements can occur between non-adjacent spatial locations, and the nonlocal dispersal captures such behaviours [63].
@@
Multiscale modelling approaches of materials can capture one or more physical phenomena at different lengths and time scales [64, 65]. Materials generally exhibit and secrete different phenomena at different lengths and time scales [65]. Due to the lack of measures and length scales in classical mechanics, the material’s microstructural features (e.g., micro-slipping and twinning), geometries, and topologies (e.g., band-gap) are unpredictable. Hierarchical and hybrid are the two types of multiscale modelling approaches studied in the literature. In the case of the hierarchical multiscale model, the material is broken down into multiple scales, and models at the different scales are run independently. In the end, coupling parameters are introduced to couple those scales. On the other hand, all different scales work together as a single package for hybrid multiscale models. The material consists of interconnected multiple subdomains over short and long ranges, and the nonlocal continuum mechanics model can capture their mechanics.
@@
Nonlocal models are capable of capturing some behaviour of biological materials such as cortical bone fracture [66], the rupture in lipid membranes [67], tumor shrinkage [68], etc. Peridynamics is one of the main theoretical and computational frameworks that unify the mechanics of discrete and continuous media [68]. Since the first papers were published, peridynamics methodology has been used to model discontinuous and long-range forces using pair-wise interactions between particles [69]. This methodology is nonlocal because it considers the long-range interaction between particles. As it deals with continuous and discrete media, it is a compelling model of biological tissue on the cell population scale. In addition, piezoelectric and other smart materials are the most commonly used medical materials. For instance, ultrasound scanners use a piezoelectric transducer; they use the converse piezoelectric effect to send out sound vibrations reflected by muscles, organs, etc. [70].
@@
Mathematical and computational models play a pivotal role in unravelling the intricate and dynamic processes within the life sciences. These models provide a systematic framework to analyze, simulate, and predict biological phenomena, aiding researchers and practitioners in understanding the underlying principles governing life at various levels of complexity. Life sciences encompass various disciplines, from molecular biology and genetics to ecology and neuroscience. Understanding the behaviour and interactions of biological systems requires integrating data and theories into mathematical and computational frameworks. These models enable the abstraction of biological processes into mathematical equations, algorithms, or simulations, allowing for a quantitative and often more precise representation of biological phenomena.
@@ 2.1 Homogeneous and heterogeneous systems
Homogeneous and heterogeneous systems are fundamental concepts in mathematical biology that play crucial roles in understanding various biological phenomena. Homogeneous systems are the baseline for understanding the fundamental dynamics of biological systems before incorporating more complex factors. For instance, patterns are observed in nature and have visible regularities. These patterns recur in different contexts and can sometimes be modelled mathematically. According to Turing [3], “when two chemical species with different diffusion rates react with each other, the spatially homogeneous state may become unstable, thereby leading to a nontrivial spatial structure”. This looks counter-intuitive as it is expected to be a uniform distribution of the chemical substances through the diffusion term. Diffusion is the process by which particles spread from high-concentration regions to low-concentration regions. Under most straightforward assumptions, Fick’s law guides this physical process of random motions for molecules. However, in some applications, other physical laws that account for relaxation processes may be required (e.g., [71]). In general, the diffusion term in RD equations spreads the concentrations throughout the spatial domain, and the linear or nonlinear reaction terms come from the mass action laws [4, 72]. Without loss of generality, we can consider a general form of RD equations with $m$-substances $\mathbf{u}(\mathbf{x},t)=(u_{1}(\mathbf{x},t),\ldots,u_{m}(\mathbf{x},t))^{T}$ at time $t>0$, and at $\mathbf{x}$ in a spatial domain $\Omega\subseteq\mathbb{R}^{n}$ as [4]:
@@
Depending on the parameters and interactions between organisms, the RD equation (1) exhibits different types of dynamic non-homogeneous solutions, such as travelling waves, wave-trains, spatio-temporal chaos, etc. In addition, reaction-diffusion systems can produce non-homogeneous stationary patterns such as spots, stripes, and target patterns [4]. In 1937, Fisher and Kolmogorov, Petrovskii, and Piskunov introduced the scalar RD equation [73]:
@@
and studied the existence of the travelling wave solutions along with the propagation of the wave speed. The equation (2) is known as the Fisher-KPP equation. Researchers have been studying and applying this Fisher-KPP model in various fields, e.g., physics, chemistry, biology, and medicine. A particular nonlinear form $g(u)=u(1-u)$ of the RD equation (2) has been taken a lot of attention in capturing different phenomena through mathematical modelling, e.g., spatial spreading of cell populations [74], wound-healing cell migration [75, 76], protein misfolding [77, 78], biological invasion [79, 80, 81], etc.
As discussed earlier, the RD equations (1) are capable of producing different types of stationary and non-stationary non-homogeneous patterns. Several mechanisms are available for getting non-stationary patterns [4]. Turing suggested a mechanism to obtain some of the non-homogeneous stationary patterns mathematically, called the Turing pattern [3]. One of the main conditions for existing Turing patterns requires at least two interacting populations for the RD equations. Therefore, the usual Turing mechanism can not be applied to the single-species local RD model (2). However, it can be applied to a single species RD model with nonlocal interactions. We first describe a nonlocal version of the Fisher-KPP RD model for a single species population (2), then move to the multi-species populations and other nonlocal interactions.